# SEAL: Towards Safe Autonomous Driving via Skill-Enabled Adversary Learning for Closed-Loop Scenario Generation

### v2

@@ V-C Metrics
For evaluating generated scenario quality, we examine the induced ego Success rate, across all tested ego methods. We derive a realism metric based on distributional measures, following MixSim [14] and other prior work [15, 33]. In particular, we utilize the Wasserstein distance (WD) over adversarial “profiles”—normalized histograms constructed from the adversary’s yaw rates, acceleration values, and out of road rates. All WD values are computed via comparison to profiles derived from the original $X_{\text{adv}}$ in $\mathbf{S}$. Furthermore, we perform a simple average over the three derived WD values to compute an overall Realism meta-metric.
@@
(a) WOMD-Normal, Real Scene
@@
Figure 3: Qualitative examples of scenario-perturbation; the ego agent is blue while the adversary is red. The original scenario from WOMD-Normal is shown in (a), while (b, c, d) demonstrate various safety-critical generation approaches.
VI Results
We report the interquartile mean (IQM) and interquartile range (IQR) over four seeds, as recommended for statistical robustness [54]. These statistical summaries are computed independently over each metric, so Success, Crash, and Out of Road may not sum to $100\%$. We also evaluate a non-reactive ego replay policy (Replay), which rolls out the original $X_{\text{ego}}$ trajectory, as well as a ReSkill [48] agent trained without any adversarial scenario generation (No Adv). Note that due to re-simulation limitations, Replay in WOMD-Normal and WOMD-SafeShift-Hard may have a nonzero failure rate.
Downstream Performance. Our closed-loop training results are summarized in Table I. SEAL-trained policies average a $\mathbf{21.5\%}$ increase in Success rate relative to the top baseline in each setting, achieving a strong balance between Crash and Out of Road rates. While a baseline-trained policy may have slightly better performance on one failure type, it is achieved by sacrificing performance against the other.
Scenario Generation Quality. To directly assess scenario generation quality, we aggregate metrics in Table II, averaged over all ego methods. Note that WOMD-Normal and WOMD-SafeShift-Hard can have non-zero WD values due to early episode terminations from ego policies’ failures. Although CAT scenes induce a lower ego Success rate than SEAL scenes, SEAL scenes exhibit the highest Realism among scenario generation approaches, a $\mathbf{35.3\%}$ improvement, contributing to SEAL-trained policies’ superior downstream performance. We also showcase qualitative examples of the tested scenario generation approaches in Figure 3, against Replay. CAT and GOOSE perturbations both result in aggressive, colliding scenarios, while the SEAL perturbation creates a more nuanced near-miss scenario, reactively avoiding a direct collision.
Ablation Studies. To further investigate how different components of SEAL affect downstream training, we perform extensive ablation studies shown in Figure 2, as well as comparing against CAT as it is a slightly stronger baseline than GOOSE. We study the effect of our learned objective function by comparing it to the heuristic, bounding box overlap approach used by CAT (Learned Obj and Heuristic Obj, respectively). Similarly, we compare our adversarial skill policy (Adv Skill Prior) with a benign prior variant (Benign Skill Prior) and a predefined trajectory following policy (TrajPred Adv). Our full SEAL approach performs the best across all evaluation settings; both the learned objective function and adversarial skill policy appear to be needed.
@@ VII Conclusion
As autonomous driving (AD) systems advance, ensuring safety remains essential. While recent safety-critical scenario generation techniques show promise, they often lack the realism, reactivity, and nuance needed to provide strong training signals for closed-loop agents. We thus introduced Skill-Enabled Adversary Learning (SEAL) as a perturbation-based safety-critical scenario generation approach, combining a learned objective function and an adversarial skill policy. In all test settings—across both real-world challenging scenarios and generated scenarios by SEAL and other SOTA methods—SEAL-trained policies achieved significantly higher success rates, with a more than 20% relative increase. Upon deeper analysis, SEAL-generated scenes contain less aggressive but more realistic adversaries, helping to explain the observed ego agent improvements. We argue that realism metrics, downstream task utility, and out-of-distribution evaluation settings are vital in assessing adversarially-perturbed scenarios.

### v3

@@ V-C Metrics
For evaluating generated scenario quality, we examine the induced ego Success rate, across all tested ego methods. We derive a realism metric based on distributional measures, following MixSim [14] and other prior work [15, 33]. In particular, we utilize the Wasserstein distance (WD) over adversarial “profiles”—normalized histograms constructed from the adversary’s yaw rates, acceleration values, and out-of-road rates. All WD values are computed via comparison to profiles derived from the original $X_{\text{adv}}$ in $\mathbf{S}$, which we average to compute an overall Realism meta-metric. We also report relative collision velocities along the contact normal, as in [25], along with head-on collision rates and severe head-on rates (where severity is defined as collision velocity exceeding 5 m/s, thereby filtering out low-speed, glancing incidents).
@@
(a) Ego Replay
(b) GOOSE [2]-Trained
(c) CAT [1]-Trained
(d) SEAL-Trained
Figure 4: Qualitative examples of driving policies; the blue ego is a learned agent while the red adversary is fixed. (a) shows the original human trajectory from WOMD-SafeShift-Hard, while (b), (c), and (d) show ego behaviors learned in different pipelines.
VI Results
We report the median and interquartile range (IQR) over four seeds, for greater statistical robustness. These statistical summaries are computed independently over each metric, so Success, Crash, and Out of Road may not sum to $100\%$. We also evaluate a non-reactive ego replay policy (Replay), which rolls out the original $X_{\text{ego}}$ trajectory, as well as a ReSkill [49] agent trained without any adversarial scenario generation (No Adv). Note that due to re-simulation limitations, Replay in WOMD-Normal and WOMD-SafeShift-Hard may have a nonzero failure rate.
Downstream Performance. Our closed-loop training results are summarized in Table I. SEAL-trained policies average a $\mathbf{21.5\%}$ increase in Success rate relative to the top baseline in each setting, achieving a strong balance between Crash and Out of Road rates. While a baseline-trained policy may have slightly better performance on one failure type, it is achieved by sacrificing performance against the other. Compared to GOOSE and CAT, SEAL training yields more realistic yaw and road compliance but less realistic acceleration, indicating stronger braking and more disciplined in-lane maneuvering to manage criticality. Despite high kinematic realism, No Adv egos crash frequently due to lack of experience in challenging scenarios and resulting poor reactivity.
We highlight qualitative examples of ego behavior in Figure 4, showcasing how different training regimes influence the execution of the same benign skills, in a scenario drawn from the WOMD-SafeShift-Hard set of real, safety-relevant scenarios. While all ego policies operate within the same offline-learned skill space, their online adaptation differs across training methods. The Replay ego depicts the ground-truth human trajectory, which merges safely into the right lane. The GOOSE-trained ego initiates the merge too early and fails to recover in time, resulting in a collision. The CAT-trained ego begins to merge later but hesitates under pressure, slows down, and is rear-ended. In contrast, the SEAL-trained ego merges with a sufficient gap while accommodating a close-following tail vehicle, resulting in a smooth and safe maneuver. These differences highlight how SEAL’s more realistic and nuanced adversarial training scenarios better prepare ego policies to navigate challenging interactions effectively.
(a) No Perturbation
@@ VI Results
Figure 5: Qualitative examples of scenario perturbation; the blue ego follows a fixed replay policy while the red adversary is modified. (a) shows the original WOMD-Normal scenario, while (b), (c), and (d) show perturbations generated by GOOSE, CAT, and SEAL, respectively.
Scenario Generation Quality. To directly assess scenario generation quality, we aggregate metrics in Table II, averaged over all ego methods. Although CAT scenarios induce a lower ego Success rate and raw head-on rate than SEAL scenarios, SEAL scenarios exhibit the highest Realism among scenario generation approaches, a $\mathbf{35.3\%}$ improvement, contributing to SEAL-trained policies’ superior downstream performance.Furthermore, SEAL’s collision velocities and severe head-on rates are far lower than baseline approaches.
We also showcase qualitative examples of the tested scenario generation approaches in Figure 5, using a fixed ego Replay policy to isolate differences in adversary behavior. CAT and GOOSE both produce aggressive trajectories that lead to collisions: CAT stops and turns directly into the ego, while GOOSE swerves across the lane and slows in the ego’s path to force a t-bone. In contrast, the SEAL adversary exhibits more nuanced behavior, slowing down to let the ego catch up, moving away at the last moment to induce a near-miss, and thus demonstrating interesting adversarial skill behavior.
Ablation Studies. To further investigate how different components of SEAL affect downstream training, we perform extensive ablation studies shown in Figure 3, as well as comparing against CAT as it is a slightly stronger baseline than GOOSE. We study the effect of our learned objective function by comparing it to the heuristic, bounding box overlap approach used by CAT (Learned Obj and Heuristic Obj, respectively). Similarly, we compare our adversarial skill policy (Adv Skill Prior) with a benign prior variant (Benign Skill Prior) and a predefined trajectory following policy (TrajPred Adv). We also compare SEAL against two additional ablations: one trained without curriculum, and another without the initial non-reactive start. Our full SEAL approach performs best across all settings; both the learned objective function and adversarial skill policy are essential, while the curriculum and non-reactive start further improve performance.
@@ VII Conclusion
As autonomous driving (AD) systems advance, ensuring safety remains essential. While recent safety-critical scenario generation techniques show promise, they often lack the realism, reactivity, and nuance needed to provide strong training signals for closed-loop agents. We thus introduced Skill-Enabled Adversary Learning (SEAL) as a perturbation-based safety-critical scenario generation approach, combining a learned objective function and an adversarial skill policy. In all test settings—across both real-world challenging scenarios and generated scenarios by SEAL and other SOTA methods—SEAL-trained policies achieved significantly higher success rates, with a more than 20% relative increase. Upon deeper analysis, SEAL-generated scenarios contain less aggressive but more realistic adversaries, helping to explain the observed ego agent improvements. We argue that realism metrics, downstream task utility, and out-of-distribution evaluation settings are vital in assessing adversarially-perturbed scenarios.
@@
ACKNOWLEDGMENT
This work was partially performed during Benjamin Stoler’s internship at Stack AV; the authors thank Stack for their mentorship. The authors additionally thank Evan Lohn for many valuable discussions throughout the design process.